# Think Smart, Act SMARL! Analyzing Probabilistic Logic Driven Safety in Multi-Agent Reinforcement Learning

### v1

@@ 2.4. ProbLog Shields
As mentioned previously, Yang et al. (2023) implement PLS using ProbLog (De Raedt et al., 2007). These programs, notated $\mathcal{T}_{(\cdot)}$, have three components: ($i$) a set of predicates (with cardinality $|\mathcal{A}|$) whose values are set to the agent’s base action distribution $\pi$, ($ii$) a set of probabilistic inputs relevant to the constraints that describe the current state, and ($iii$) a set of sentences that describe the agent’s constraints, specifically including a definition for a predicate ‘safe_next’ (the safety of the next state under $\pi$). Here, ‘safety’ is used as a general term for any probabilistic constraint satisfaction.
@@
8unsafe_next :- action(stag), sensor(stag_diff).
9unsafe_next :- action(hare), sensor(hare_diff).
10safe_next :- \+unsafe_next.
@@
An exemplary shield (above, Shield 1) that was used in the Stag-Hunt experiments for constraining the agents towards the mixed Nash equilibrium is described as such: The values of the ‘action($\cdot$)’ predicates are inferred from $\pi$. The ‘sensor($\cdot$)’ predicates are set to the normalized absolute difference between the precomputed mixed Nash strategy and the mean of the agent’s historical actions. The predicate ‘safe_next’ is inferred via probabilistic logic semantics and determine the safety of the current state.
3. Methods
In this section, we introduce two methods that integrate PLS into MARL settings. First, in Section 3.1, we describe Probabilistic Logic TD Learning (PLTD), which mirrors shielded policy gradient methods within temporal difference learning by incorporating safety constraints. Following this, we present Probabilistic Logic SMARL in Section 3.2, which applies the concept of probabilistic shielding to multi-agent environments. In SMARL, each agent operates independently with its own shield, allowing flexibility in experimentation and parameter sharing.
3.1. Probabilistic Logic Shielded Temporal Difference Learning (PLTD)
Yang et al. (2023) define their shielded algorithm, PLPG, with respect only to policy gradient methods, specifically PPO (Schulman et al., 2017). Thus, in order to use it with DQN (Mnih et al., 2013), SARSA (Zhao et al., 2016) (or other deep TD-learning algorithm), we must first define how an agent uses a ProbLog shield to learn a shielded policy during exploration and evaluation. An important consideration for extending PLS to TD-learning is whether the algorithm is on-policy or off-policy. Yang et al. (2023) note that for off-policy algorithms to converge, both the exploration and learned policies must cover the same state-action space. Since in PLS, the agent optimizes $\pi$ with $\pi^{+}$, one must be careful about this assumption. We propose both on- and off-policy algorithms that incorporate safety constraints into Q-learning.
3.1.1. Objective Function
The full specifications of PLPG include a shielded policy gradient $[\nabla_{\theta}\log P_{\pi^{+}}(\mathtt{safe}|s)]$ and a safety gradient penalty $[-\log P_{\pi^{+}}(\mathtt{safe}|s)]$. Since TD-methods are not policy gradient methods, we must rely on the latter to introduce information about safety constraints into the loss function.
Let $\mathcal{D}$ be the distribution of $d=\langle s^{t},a^{t},r^{t},s^{t+1},a^{t+1}\rangle$ tuples extracted from a history buffer of the agent interacting with the MDP and $s^{t}$, $a^{t}$, $r^{t}$ be the state, taken action, and reward at time $t$. Let the Q-value approximator $Q_{\theta}(s,a)$ be parameterized by some parameters $\theta$. Then, the off-policy loss (Q-learning based) and on-policy loss (SARSA-based) are augmented using this penalty term below to create Probabilistic Logic TD Learning (PLTD).
Definition 1 (Probabilistic Logic TD Learning, PLTD).
The PLTD minimization objective is:
3.2. Probabilistic Logic SMARL
In this paper, at least one of the agents of an independent MARL algorithm uses PLS. This is called Probabilistic Logic Shielded MARL. Two SMARL algorithm variations (using neural network approximators) of Independent Q-Learning (IQL) (Tan, 1993; Albrecht et al., 2023) and three of Independent PPO (IPPO) (De Witt et al., 2020) based on the level of parameter sharing between agents can be found in Table 1.
Algorithm Acronym Base Algo. Shared Parameters
Shielded IQL SIQL PLTD None
Shielded PSQL SPSQL PLTD Policy Network
Shielded IPPO SIPPO PLPG None
Shielded CSPPO SCSPPO PLPG Critic Network
Shielded PSPPO SPSPPO PLPG Critic & Actor Networks
Table 1. Shielded MARL algorithms used in this paper.
All SMARL algorithms that are discussed herein have agents being relatively independently shielded. Each agent also may have their own independent policies and independent parameter updates. This allows for maximal flexibility during experimentation (for example, testing different degrees of parameter sharing or partially-shielded populations). A schematic diagram of the parallel interaction of the agents with respect to some environment can be seen in Figure 2. The agent-environment-cycle interaction scheme, used often in turn-based games (Terry et al., 2021) can similarly be easily extended to the SMARL framework by having each agent interact with their own shield before sampling an action.
Figure 2. SMARL interaction schematic where all agents take actions in parallel.
In Figure 2, each agent $p_{i}$, indexed with $i\in\{1,...,n\}$, observes a possibly subjective current state $s_{i}^{t}$ and sends a policy $\pi_{i}(s^{t}_{i})$ and the current state (potentially modified) $s_{i}^{t}$ to their shield $\mathcal{T}_{i}$. Each $\mathcal{T}_{i}$ is a ProbLog program from which we can query various safety related probability measures to compute a safe policy $\pi_{i}^{+}(s_{i}^{t})$. An action $a_{i}^{t}$ is sampled from the safe policy using some exploration or exploitation method. Finally, the joint set of actions ${a}^{t}$ is sent to the environment $Env$. After updating its state, $Env$ sends new observations (the tuple $\mathbf{s}^{t+1}$) and rewards (${r}^{t+1}$) to the agents, and the cycle continues with $t\leftarrow t+1$ until reaching a terminal state.
4. Results
The results for probabilistic logic SMARL for two agents in the various game-theoretic environments described earlier in Section 2.1 are presented below. For the sake of clarity, details are moved to the supplementary material: the payoff matrices/reward structure for each game can be found in Appendix A; shield constructions and programs for each experiment can be found in Appendix B; to validate the efficacy of PLTD with respect to standard DQN, a single-agent experiment is described in Appendix C; and all hyperparameters can be found in Appendix D. Our code can be found at https://github.com/satchitchatterji/ShieldedMARLThesis (Chatterji, 2024).
4.1. Stag-Hunt
Figure 3. Training, evaluation and safety results for Stag-Hunt for PLPG-based agents. Results are smoothed using a rolling average with window size 50. The lines represent the mean and the shadow the standard deviation over 5 seeds.
@@ 4.1. Stag-Hunt
Stag-Hunt was used to test the ability of PLS to guide agents towards pre-selected behaviors in games with multiple Nash equilibria. Figure 3 shows the results of PPO-based agents in the repeated Stag-Hunt game. IPPO serves as the baseline due to the environment’s simplicity, while SIPPO agents are equipped with a cooperation-inducing shield (Shield 2 $\mathcal{T}_{pure}$). The IPPO agents converge to the non-cooperative Nash equilibrium, consistently playing Hare with a reward of 2 per step. In contrast, agents using $\mathcal{T}_{pure}$ quickly and reliably play Stag, earning a higher reward of 4. Unshielded agents fail to reach the mixed strategy due to the instability of PPO’s stochastic policies – small deviations push agents toward pure strategies, often resulting in the non-cooperative equilibrium. However, agents using $\mathcal{T}_{mixed}$ (Shield 3) successfully adopt a mixed strategy, as indicated by the expected rewards and safety measures, despite high variability, likely caused by the instability of the normative strategy relative to pure Nash equilibria.
4.2. Centipede Game
DQN, $\epsilon$-greedy
DQN, softmax
Figure 4. Training, evaluation and safety results for Centipede for PLPG- and PLTD-based agents. The PLTD-based agents have exploration policies $\epsilon$-greedy (evaluations executed with a greedy policy) and softmax (evaluations executed with a softmax policy). The lines represent the mean and the shadow the standard deviation over 5 seeds for PLPG agents and 3 seeds for PLTD agents. Results are smoothed using a rolling average with window size 50.
@@ 4.2. Centipede Game
Centipede Game was used to test how well PLS can guide agents to cooperate in order to achieve higher collective rewards, even in scenarios where the dominant strategy suggests defection – a shield was constructed (Shield 4) to encourage agents to push the game further by constraining early defection behavior. Figure 4 show results for pairs of MARL agents learning to play the Centipede game. IPPO agents occasionally learn to play Continue through all 50 steps within 500 episodes, and it is likely they would all converge if trained further. In contrast, we see the SIPPO agents playing the full game from the start.
@@
For DQN-based agents, all SIQL variations cooperate by playing Continue from the beginning. In contrast, $\epsilon$-greedy agents fail to progress far, as early on in exploration, $\epsilon\approx 1$, meaning there is a uniform distribution over the actions, and only a 25% probability of the agents going to only the second stage of the game. The probability of reaching some early stage $n$ is thus $\approx 0.25^{n-1}$. This discourages exploration, as stopping early provides a small but consistent reward, aligning with the Nash equilibrium (i.e. the subgame-perfect equilibrium). Unshielded softmax agents perform slightly better, but with high variance – some complete the game while most exit early. Shielding enables all agents to explore much further into the game.
@@
4.3. Extended Public Goods Game
@@
Figure 5. Training, evaluation and safety results for the final 10 episodes for the 2-player Extended Public Goods Game for PLPG-based agents (mean results over 5 seeds). The error bars show standard deviation. For a fuller analysis including training curves, please refer to Appendix E.
EPGG was used to test the robustness of PLS in promoting cooperation in managing uncertainty in environment with stochastic payoffs. A shield (Shield 5) was designed to enable the agents to play the expected Nash equilibrium (as if it was playing with payoffs determined by an estimate $\hat{\mu}$) and not the instantaneous one (with the equilibrium determined by $f_{t}$). Figure 5 shows the results of agents playing the EPGG with $f_{t}\sim\mathcal{N}(\mu,1)$ where $\mu\in\{0.5,1.5,2.5,5.0\}$.
We see that in all cases, the IPPO agents (baseline) learn to play the Nash solution to the instantaneous EPGG – this is apparent from the reward curves and the mean cooperation, where it is better to defect when $f_{t}<1$ and cooperate when $f_{t}>2$, with mixed incentives when $21<f_{t}<$ 2. The SIPPO agents quickly learn the expected Nash erring on the side of cooperation in mixed scenarios, as directed by the shield. This leads to more prosocial behavior – defection when $\hat{\mu}<1$ and cooperation when $\hat{\mu}>1$. We also see that the further away $\mu$ is from a mixed incentive scenario, the lower the variance.
For $\mu=0.5$, SIPPO agents show less variation in reward and behavior than IPPO, converging quickly to the Nash strategy (non-cooperation) but with smaller rewards. This is due to the fact that when $f_{t}\sim\mathcal{N}(0.5,1)$, there is a $\approx 30$% chance of the value being greater than 1, which, through cooperation, the agents can attain higher instantaneous rewards than if they play the expected strategy. Since the shield is designed to help learn the expected strategy rather than the instantaneous strategy, the fact that the agents obey the shield rather than the reward is a positive result. Similarly, at $\mu=5$, SIPPO agents again exhibit lower variance, always cooperating. As $\mu$ is interpolated between 0.5 and 5, the optimal strategy flips from defect to cooperate. We see that for the shielded agents, the uncertainty whether to play this strategy is maximum when $\mu=1$, where the agents play each action approximately equally, which also leads to higher rewards than the baseline.
Except for $\mu=0.5$, the rewards for the SIPPO agents are either equal to or exceed the unshielded ones, and their cooperative or defective nature are more predictable.
4.4. Markov Stag-Hunt
The Markov Stag-Hunt experiments investigate ($i$) the ability of PLS to promote cooperation in dynamic settings, ($ii$) the effect of shields with different constraining properties in the same environment (4.4.1), and ($iii$) the effect of shields when a part of the population is shielded (with and without parameter-sharing, 4.4.2).
4.4.1. Variation in Shield Strength
Figure 6. Behavioural and training results during training for Markov Stag-Hunt for PLPG-based agents with no shield or two different shields, $\mathcal{T}_{weak}$ or $\mathcal{T}_{strong}$. The lines represent the mean and the shadow the standard deviation over 3 seeds. Results are smoothed using a rolling average with window size 50.
Figure 7. Behavioural and training results for Markov Stag-Hunt for PLPG-based agents with or without parameter sharing during training, where one of the agents is shielded with $\mathcal{T}_{strong}$. The solid lines represent the mean and the shadow the standard deviation over 3 seeds. Results are smoothed using a rolling average with window size 50.
Two shields, $\mathcal{T}_{strong}$ (Shield 7) and $\mathcal{T}_{weak}$ (Shield 8), were constructed such that $\mathcal{T}_{strong}\subset\mathcal{T}_{weak}$ in terms of their safety constraints. Broadly, $\mathcal{T}_{weak}$ constrains the agents’ behaviors to strongly cooperate when they are near the stag, and $\mathcal{T}_{strong}$ constrains them regardless of where they are placed on the grid. Figure 6 presents the agents’ behavior. In this experiment, analyzing reward or safety alone does not fully capture cooperative behavior, so the key metrics are: (6(a)) plant harvests per episode, (6(b)) successful stag hunts with another agent, and (6(c)) attempts to hunt a stag alone. Additional graphs show training (6(d)) and evaluation (6(e)) rewards, and mean safety for shielded agents (6(f)), verifying their increasing rewards and safety over time.

### v2

@@ 3.3 Shield Construction using ProbLog
A shield in PLS is defined using a ProbLog program [9], denoted $\mathcal{T}$. Each consists of three components: ($i$) an annotated disjunction $\Pi_{s}$ of predicates whose values are set to a base action distribution $\pi$, ($ii$) a set of constraint-related inputs describing the current state $\mathbf{H}_{s}$, and ($iii$) a set of sentences $\mathcal{KB}$ (for knowledge base) that specify the agent’s constraints, including a definition for a predicate ‘safe_next’ (the predicted safety of the next state under $\pi$). Again, ‘safety’ broadly refers to any probabilistic constraint satisfaction goal (ref. Sec 2.1).
An exemplary shield (Shield 1) for constraining agent behavior towards the mixed Nash equilibrium in Stag-Hunt games is described as such: The probabilistic valuations of the ‘action($\cdot$)’ predicates are inferred from $\pi$. The ‘sensor($\cdot$)’ predicates are set to the normalized absolute difference between the precomputed mixed Nash strategy and the mean of the agent’s historical actions. The predicate ‘safe_next’ is inferred via PL semantics [8] and determines the safety of the next state under $\pi$ – safer states are those where the stated normalized absolute differences are low. For brevity, we describe subsequent shields in-line rather than displaying its full ProbLog source.
@@
8unsafe_next :- action(stag),
9 sensor(stag_diff).
10unsafe_next :- action(hare),
11 sensor(hare_diff).
12safe_next :- \+unsafe_next.
@@
3.3.1 Deterministic Shields
Using PLS, agents may be made to be deterministic in their exploration and deployment which may be useful when training agents in environments requiring hard constraints. This can be formulated as the following (proven in the supplementary):
Proposition 3.5.
For any game there exists a shield $\mathcal{T}$ such that the resulting shielded agent deterministically selects a single action $a\in\mathcal{A}$ for each state $s\in S$ if the shielded policy $\pi^{+}$ computes all other actions $a^{\prime}\in\mathcal{A}\setminus\{a\}$ as perfectly unsafe.
4 Results
We organize our experiments around key multi-agent phenomena relevant to several open MARL challenges. Rather than focusing on games individually, each section centers on a particular strategic challenge (e.g., equilibrium selection, social dilemmas) and uses one or more games as illustrative testbeds. For the sake of clarity, details are moved to the supplementary material, such as implementation details, shield programs for each experiment, a validation of PLTD in a single-agent experiment, all hyperparameters and training curves. All displayed results are averaged over 5 random seeds to account for variability in training. All reported errors correspond to standard deviations across runs and respective episodes/steps. Our code can be found at https://github.com/satchitchatterji/ShieldedMARLThesis.
4.1 Coordination and Equilibrium Selection (Stag-Hunt)
We use Stag-Hunt to test the ability of PLS to guide agents towards normative behaviors in games with multiple Nash equilibria. IPPO serves as the baseline, while SIPPO agents are equipped with a shield constraining the agents towards either the pure ($\mathcal{T}_{pure}$) or mixed Nash equilibrium ($\mathcal{T}_{mixed}$). $\mathcal{T}_{pure}$ simply tells the agent to not play $Hare$, making it a deterministic shield according to Proposition 3.5. $\mathcal{T}_{mixed}$, used as an example in Section 3.3, defines safety as the absolute difference between the precomputed mixed Nash strategy and the mean of the agent’s historical actions.
Table 1: Results for Stag-Hunt with PPO-based agents over the last 50 training episodes. $\bar{r}$ is the mean return and cooperation is defined as ${t_{max}^{-1}}\sum_{t=0}^{t_{max}}\mathbf{P}_{\mathcal{T}_{pure}}(\texttt{safe}|s^{t})$.
Algorithm $\bar{r}$ (train) $\bar{r}$ (eval) cooperation
IPPO 1.99$\pm$0.03 1.99$\pm$0.02 0.01$\pm$0.01
SIPPO ($\mathcal{T}_{pure}$) 4.00$\pm$0.00 4.00$\pm$0.00 1.00$\pm$0.00
SIPPO ($\mathcal{T}_{mixed}$) 2.57$\pm$0.48 2.63$\pm$0.43 0.58$\pm$0.08
@@ 4.1 Coordination and Equilibrium Selection (Stag-Hunt)
Table 1 shows the results this setting. The IPPO agents converge to the non-cooperative Nash equilibrium, consistently playing Hare with a reward of 2 per step. These unshielded agents fail to reach the mixed Nash equilibrium likely due to the instability of PPO’s stochastic policies – small deviations push agents toward pure strategies, often resulting in the non-cooperative equilibrium. In contrast, agents using $\mathcal{T}_{pure}$ quickly and reliably play Stag, earning a higher reward of 4. Agents using $\mathcal{T}_{mixed}$ successfully adopt a mixed strategy, as indicated by the expected rewards and cooperation measures, despite high variability, likely caused by the instability of the normative strategy relative to those under the two pure Nash equilibria.
4.2 Temporal Coordination (Centipede)
@@
We use Centipede to test how well SMARL guides agents to cooperate to achieve large collective long-term rewards, even in scenarios where the dominant strategy suggests defection. A shield was constructed ($\mathcal{T}_{continue}$) to encourage agents to play the game by constraining against early defection. Table 2 shows results for pairs of MARL agents learning to play the Centipede game. IPPO agents occasionally learn to play Continue through all $t_{max}=50$ steps within 500 episodes. In contrast, we see the SIPPO agents playing the full game from the start of training.
Table 2: Results for Centipede with PPO- and DQN-based agents over the last 50 training episodes. Safety is defined as $\frac{1}{T}\sum_{t=0}^{T}\mathbf{P}_{\mathcal{T}_{cent}}(\texttt{safe}|s^{t})$ and $R_{\text{ep}}=\sum_{t=0}^{{t_{max}}}r^{t}$.
Algorithm $R_{\text{ep}}$ (train) $R_{\text{ep}}$ (eval) Safety
IPPO 42.35$\pm$36.62 42.83$\pm$35.81 0.83$\pm$0.23
SIPPO 100.50$\pm$0.00 100.50$\pm$0.00 1.00$\pm$0.00
IQL ($\epsilon$-greedy) 34.62$\pm$46.59 34.70$\pm$46.53 0.68$\pm$0.23
SIQL ($\epsilon$-greedy) 100.50$\pm$0.00 100.50$\pm$0.00 1.00$\pm$0.00
IQL (softmax) 1.73$\pm$1.01 30.10$\pm$38.61 0.73$\pm$0.21
SIQL (softmax) 100.50$\pm$0.00 100.50$\pm$0.00 1.00$\pm$0.00
Figure 3: Results for the final 10 episodes of the 5-player EPGG for SIPPO, SCSPPO, and SACSPPO agents with varying $\mu$ and ratios of shielded agents.
@@
For DQN-based agents, $\epsilon$-greedy and softmax exploration policies were tested. Unshielded $\epsilon$-greedy agents fail to progress far, as early on in exploration, $\epsilon\approx 1$ ($\pi$ is uniform), and thus a 25% probability of the agents going to the second stage of the game. The probability of reaching some early stage $\ell$ is $\approx 0.25^{\ell-1}$. This discourages exploration, as stopping early provides a small but consistent reward, aligning with the Nash/subgame-perfect equilibrium. Unshielded softmax agents perform slightly better, but with high variance – some agent pairs cooperate while most exit early. This highlights the influence of the choice of exploration strategy, and how using shield can guide desirable behavior – all SIQL variations learn to play Continue consistently.
@@
4.3 Social Dilemmas in Stochastic, Mixed-Motive Environments (Extended Public Goods Game)
We use EPGG to investigate how PLS can promote cooperation in environments with stochastic social incentives. Our experiments address two central questions: ($i$) can agents learn to adopt normative strategies that reflect long-term expectations rather than short-term payoffs? and ($ii$) how does the effectiveness of shielding scale when only a subset of agents is shielded, especially under different degrees of parameter sharing? In the first experiment, we assess how shielding can steer agents toward the expected Nash equilibrium of the EPGG. In the second, we evaluate how the number of shielded agents and the degree of parameter sharing affect emergent cooperation and reward across a 5-agent population.
4.3.1 Learning Long-Term Normative Behavior
We begin with a two-player setting to isolate the effect of shielding on equilibrium selection under payoff uncertainty. This controlled setup enables a clear comparison between shielded agents following the expected Nash equilibrium and unshielded agents reacting to stochastic payoffs. Limiting the population removes confounding factors such as social influence, and allows for straightforward game-theoretic analysis, providing a clean baseline before scaling to larger populations.
To test whether PLS can guide agents toward long-term cooperation, we designed a shield ($\mathcal{T}_{EPGG}$) that directs agents to follow the expected Nash equilibrium, based on an empirical estimate $\hat{\mu}$ of the stochastic payoff multiplier $f_{t}\sim\mathcal{N}(\mu,1)$. This contrasts with the unshielded baseline, where agents maximize the instantaneous payoff.
Figure 4 reports results for $\mu\in\{0.5,1,1.5,2.5,5.0\}$. Unshielded IPPO agents reliably converge to the instantaneous equilibrium – defecting when $f_{t}<1$ and cooperating when $f_{t}>2$. In contrast, SIPPO agents quickly converge to the desired behavior, leading to more prosocial behavior under uncertainty.
We observe that SIPPO agents exhibit lower reward variance and more stable cooperation compared to the baseline, especially in mixed-motive cases ($21<\mu<$ 2). For $\mu=0.5$, the shielded agents correctly learn to defect, but earn slightly less due to foregoing occasional cooperative gains from high $f_{t}$. Overall, shielded agents match or exceed the baseline in reward and demonstrate more consistent cooperation.
@@ 4.3.1 Learning Long-Term Normative Behavior
Figure 4: Results for the final 10 episodes for the two-player Extended Public Goods Game for PLPG-based agents.
4.3.2 Partial Shielding and Parameter Sharing
To focus on how shield coverage and parameter sharing influence group behavior, we use a simplified shield that always selects the cooperative action. Unlike the previous experiment, this avoids computing expected equilibria and isolates the effect of shielding structure. Importantly, even though the shield is deterministic, shielded agents still receive gradient-based learning signals. This allows safety information to influence shared parameters, enabling unshielded agents to adopt cooperative behavior indirectly.
We run EPGG with $n=5$ agents, where a subset $k\in\{0,\dots,5\}$ of agents are shielded with a deterministic policy that always cooperates (reported as a shielded ratio=$k/n$). We compare SIPPO, SCSPPO, and SACSPPO to test how different levels of parameter sharing modulate the influence of the shielded agents on the unshielded population.
Figure 3 shows grouped bar charts for $\mu\in\{0.5,1,1.5,2.5,5,7.5\}$. Results show a clear trend: increasing the number of shielded agents leads to higher overall cooperation and greater episodic reward in the mixed-motive range ($51<\mu<$ 5). Furthermore, this effect is amplified by parameter sharing: SACSPPO consistently achieves the same or higher cooperation than SCSPPO and SIPPO at the same shielded ratio. In fully cooperative settings ($\mu=7.5$), all configurations converge to high cooperation, but SACSPPO still exhibits the lowest variance. In competitive settings ($\mu=0.5$), cooperation remains low for the baseline algorithms, as expected, increasing with the number of shielded agents – this results in an increasingly lower mean reward with the benefit of the agents instead cooperating. Notably, in the competitive regime, SACSPPO exhibits the highest level of cooperation among the methods, indicating that shared learning with shielded agents can drive prosocial behavior in unshielded agents – even when it reduces their individual rewards.
These results demonstrate that probabilistic shields can influence unshielded agents via shared parameters, serving as an effective conduit for prosocial behavior in multi-agent social dilemmas.
4.4 Spatio-Temporal Coordination with Imperfect Information (Markov Stag-Hunt)
We use the Markov Stag-Hunt environment to investigate how PLS in SMARL shapes coordination in sequential, stochastic, and partially observable multi-agent settings. This environment extends the normal-form Stag-Hunt to a spatial domain where agents must learn when and how to cooperate over time [24]. Three experiments are conducted: In Section 4.4.1, we first study the effect of shield constraint strength on agent behavior. Then, in Section 4.4.2, we evaluate how cooperation scales with population size across a range of MARL algorithms. Finally, Section 4.4.3 explores how cooperation emerges when only a subset of agents is shielded, highlighting the role of parameter sharing in diffusing safety signals across the population.
4.4.1 Variation in Shield Strength
Table 3: Results for 2-player Markov Stag-Hunt with shielded and unshielded PPO-based agents. Columns show: total plants harvested, stags killed, penalties from solo hunts, episodic reward ($R_{\text{ep}}=\sum_{t=0}^{t_{max}}r^{t}$), and mean episodic safety (${t_{max}^{-1}}\sum_{t=0}^{t_{max}}\mathbf{P}_{\mathcal{T}_{(\cdot)}}(\texttt{safe}|s^{t})$).